P5.4 Consolidation Working Paper One Year After Acquisition, Bargain Purchase (see re- lated P4.5) On January 1, 2019, Paxon Corporation acquired 90 percent of the outstanding com mon stock of Saxon Company for S1,520 million cash. The fair value of the 10 percent noncontrolling interest in Saxon was estimated to be $180 million at the date of acquisition. Paxon uses the complete equity method to report its investment. The trial balances of Paxon and Saxon at December 31, 2019 appear below Dr (Cr) In millions) Paxon Saxon Cash and receivables S 3,370 800 Equity method investments Investment in Saxon 1,942.2 650 3,600 (2,020) (5,000) 300 1,150 (1,200) (450) Buildings and equipment, net Current liabilities Long-term debt Common stock, par value. Additional paid-in capital. Retained eamings, January 1 (1,200) (2,410) Sales revenue. Equity in net income of Saxon Gain on acquisition Gain on sale ot securities Cost of goods sold Depreciation and amortization expense Interest expense Other operating expenses Totals (30,000) 10,000) (412.2) (100) (10) 8,000 300 25 1,600 2,770 Several of Saxons assets had fair values different from their book values at the acquisition date, as (in mililions) Fair value Book value $(100) 245 Inventory (FIFO) (sold in 2019) Equity method investments (sold in 2019) Buildings and equipment, net (20 years, straight-line) 300 In addition, Saxon had a 5-year life, straight-line ly unrecorded identifiable intangible assets valued at $110 million, with

Required:

a) Prepare a schedule computing the gain on acquisition.

b) Prepare a schedule calculating the equity in net income of Saxon for 2019, reported on Paxon's books, and the noncontrolling interest in net income for 2019, to be reported on the consolidated income statement for 2019.

c) Prepare a working paper to consolidate the trial balances of Paxon and Saxon at December 31, 2019.
Answer #1

Compute Gain on acquisition as follows:

Particulars Amount

Investment in Saxon

Less: Share capital (500*90%)

Less: Pre-retained earning (845*90%)

Less: Dividend of the previous year (100*90%)

1520

(450)

(760.50)

(90.00)

Goodwill 219.50

Net Income :

Investment Value

Less: Purchase Acquisition

Less: Goodwill

1942.20

(1520)

(219.50)

Net Income 202.70

Non Controlling Interest Value

Less: Share Capital

Less: Pre-Retained Earning

1912.50/90*10% = 212.50

(500*10%) = (50.00)

(845*10%) = (84.50)

Net Income 78.00